OpenStudy (anonymous):

help!! Limx->2 (((√5x)-3)/(x))

OpenStudy (lgbasallote):

\[\huge \lim_{x \rightarrow 2} \left(\frac{\sqrt{5x} - 3}{x}\right)\] is that the question?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so substitute 2..

OpenStudy (lgbasallote):

\[\huge \implies \frac{\sqrt{5(2)} - 3}{2}\]

OpenStudy (lgbasallote):

then simplify

OpenStudy (lgbasallote):

that would be your answer

OpenStudy (anonymous):

\[(\sqrt{10}-3)\div 2\] so thats it?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

how would i graph this?

OpenStudy (lgbasallote):

i have no idea. graphing is my weakest point. sorry